# EmoAttack: Utilizing Emotional Voice Conversion for Speech Backdoor Attacks on Deep Speech Classification Models

### v1

@@ Abstract
Deep speech classification tasks, mainly including keyword spotting and speaker verification, play a crucial role in speech-based human-computer interaction. Recently, the security of these technologies has been demonstrated to be vulnerable to backdoor attacks. Specifically speaking, speech samples are attacked by noisy disruption and component modification in present triggers. We suggest that speech backdoor attacks can strategically focus on emotion, a higher-level subjective perceptual attribute inherent in speech. Furthermore, we proposed that emotional voice conversion technology can serve as the speech backdoor attack trigger, and the method is called EmoAttack. Based on this, we conducted attack experiments on two speech classification tasks, showcasing that EmoAttack method owns impactful trigger effectiveness and its remarkable attack success rate and accuracy variance. Additionally, the ablation experiments found that speech with intensive emotion is more suitable to be targeted for attacks.
@@ I Introduction
Speech classification is vital in areas like autonomous driving, smart healthcare, and speaker authentication. It requires extensive data, numerous trainable parameters, and costly resources. To reduce costs and resource demands, some developers outsource data or model training to third parties. Research shows that using third-party platforms for deep neural networks(DNNs) training can introduce security risks[1], such as data or code poisoning[2, 3], where attackers embed backdoors into models. These backdoors prompt models to produce incorrect classifications when a specific trigger hides in the inputted samples, making speech classification models vulnerable to such attacks.
Backdoor attacks have been explored earlier in the area of image and text classification. For example, BadNets by Gu et al.[4] and specific sentence by Dai et al.[5], both involve inserting poisoned samples that cause incorrect model outputs when triggered. These attacks, known as poisoning-label attacks, often modify images using masking, embedding, or color adjustments. However, such techniques may not be effective for speech data. Research shows that speech and image triggers differ due to their distinct physical properties[6, 7]. Some speech-trigger methods mimic image-based approaches by inserting noise or specific sounds into speech[8, 9, 10, 11, 12, 13, 14, 15], but these are easily often detectable by humans. To enhance stealthiness, attackers now aim to modify speech components without disrupting them. Ye et al.[16] introduced a voice conversion (VC) trigger that alters timbre and associates it with a target label. Cai et al.[17] proposed using pitch and timbre as joint triggers for speech backdoor attacks.
@@
  • We compared the effectiveness of attacks targeting different emotions. The greater the emotional difference before and after the trigger, the more effective the attack.
@@ III-B Adversary’s Goals
The attacker’s goals are stealthiness, effectiveness, and robustness. Stealthiness means backdoor attacks must evade human and machine detection, with poisoned utterances resembling normal utterances. Effectiveness demands high attack success with minimal poisoning in tests. However, high success rates often require many poisoned samples, reducing stealthiness. Robustness ensures attacks resist simple detection and remain effective against adaptive defenses and in real-world conditions.
@@ III-C Proposed EmoAttack Pipeline
Poisoned Samples Generation. As shown in Figure 2, the EmoAttack aims to associate the emotional categories of utterances with the target label. It’s noted that the majority of samples are classified as neutral emotions by a SER model. We denote a speech sample $x$ with emotion $e$ and label $y$ as $(x,y,e)$. The $N$ samples in the clean dataset $D=\{(x_{i},y_{i},e_{i})\}$ are firstly divided into two parts: clean training set $D_{t}$, and clean test set $D_{c}$, noting that $D=D_{t}+D_{c}$ and $e_{i}$ denoted the emotion of $i$-th sample. After this, we randomly separated a subset from $D_{t}$ that owned the mostly same emotion, denoted as $D_{tq}$. Based on the top amount of emotion category, we applied the EVC trigger to each sample in $D_{tq}$ and modified their labels to the target label $y_{t}$, resulting in the set $D_{tp}=\{(x_{p}=EVC(x_{i},e_{t}),y_{t})\}$, where the $e_{t}$ denoted another target emotion category.
@@ IV-C Ablation Study
Detection Categories by SER Model. When selecting the samples to be attacked, we used the SER model to determine the maximum voting category of the clean dataset and selected the samples to be attacked from the data in that category. We found that not doing so may result in unstable or decreased ASR. Therefore, the better attack strategy is to attack the neutral speech.
@@ V Conclusion
This paper analyzes the differences between backdoor attacks in the domains of images and speech. We proposed EmoAttack, a backdoor attack method based on emotional voice conversion. This method preserves the linguistic content and timbre characteristics of speech while modifying a higher-level attribute of speech: emotion. After EmoAttack training, the emotional utterances can lead the victim model to produce wrong predictions. We conducted backdoor attack experiments on two speech classification tasks. The experimental results demonstrate excellent attack effectiveness of the EmoAttack. Additionally, we verified that different emotions as target labels result in varying efficiency of the trigger, the intense emotions gain better results. The proposed method aims to provide insights into backdoor attacks in the speech domain.

### v2

@@ Abstract
Deep speech classification tasks, mainly including keyword spotting and speaker verification, play a crucial role in speech-based human-computer interaction. Recently, the security of these technologies has been demonstrated to be vulnerable to backdoor attacks. Explicitly speaking, speech samples are attacked by noisy disruption and component modification in present triggers. We suggest that speech backdoor attacks can strategically focus on emotion, a higher-level subjective perceptual attribute inherent in speech. Furthermore, we proposed that emotional voice conversion technology can serve as the speech backdoor attack trigger, and the method is called EmoAttack. Based on this, we conducted attack experiments on two speech classification tasks, showcasing that EmoAttack method owns impactful trigger effectiveness and its remarkable attack success rate and accuracy variance. Additionally, the ablation experiments found that speech with intensive emotion is more suitable to be targeted for attacks.
@@ I Introduction
Speech classification is vital in areas like autonomous driving, smart healthcare, and speaker authentication. It requires extensive data, numerous trainable parameters, and costly resources. Some developers outsource data or model training to third parties to reduce costs and resource demands. Research shows that using third-party platforms for deep neural networks(DNNs) training can introduce security risks[1], such as data or code poisoning[2, 3], where attackers embed backdoors into models. These backdoors prompt models to produce incorrect classifications when a specific trigger hides in the inputted samples, making speech classification models vulnerable to such attacks.
Backdoor attacks have been explored earlier in the area of image and text classification. For example, BadNets by Gu et al.[4] and sentences attack by Dai et al.[5] both involve inserting poisoned samples that cause incorrect model outputs when triggered. These attacks, known as poisoning-label attacks, often modify images using masking, embedding, or color adjustments. However, such techniques may not be effective for speech data. Research shows that speech and image triggers differ due to their distinct physical properties[6, 7]. Some speech-trigger methods mimic image-based approaches by inserting noise or specific sounds into speech[8, 9, 10, 11, 12, 13, 14, 15], but these are easily detectable by human ears. To enhance stealthiness, attackers now aim to modify speech components without disrupting them. Ye et al.[16] introduced a voice conversion (VC) trigger that alters timbre and associates it with a target label. Cai et al.[17] proposed using pitch and timbre as joint triggers for speech backdoor attacks.
@@
  • We compared the effectiveness of attacks targeting different emotions. The more significant the emotional difference before and after the trigger, the more influential the attack.
@@ III-B Adversary’s Goals
The attacker’s goals are stealthiness, effectiveness, and robustness. Stealthiness means backdoor attacks must evade human and machine detection, with poisoned utterances resembling normal utterances. Effectiveness demands high attack success with minimal poisoning in tests. However, high success rates often require many poisoned samples, reducing stealthiness. Robustness ensures attacks resist simple detection and remain effective against adaptive defences in real-world conditions.
@@ III-C Proposed EmoAttack Pipeline
Poisoned Samples Generation. As shown in Figure 2, the EmoAttack aims to associate the emotional categories of utterances with the target label. It’s noted that most samples are classified as neutral emotions by a SER model. We denote a speech sample $x$ with emotion $e$ and label $y$ as $(x,y,e)$. The $N$ samples in the clean dataset $D=\{(x_{i},y_{i},e_{i})\}$ are firstly divided into two parts: clean training set $D_{t}$, and clean test set $D_{c}$, noting that $D=D_{t}+D_{c}$ and $e_{i}$ denoted the emotion of $i$-th sample. After this, we randomly separated a subset from $D_{t}$ that owned the mostly same emotion, denoted as $D_{tq}$. Based on the top amount of emotion category, we applied the EVC trigger to each sample in $D_{tq}$. We modified their labels to the target label $y_{t}$, resulting in the set $D_{tp}=\{\overline{x_{i}}=EVC(x_{i},e_{t}),y_{t})\}$, where the $e_{t}$ denoted another target emotion category.
@@ IV-C Ablation Study
Detection Categories by SER Model. When selecting the samples to be attacked, we used the SER model to determine the maximum voting category of the clean dataset. We selected the samples to be attacked from the data in that category. We found that not doing so may result in unstable or decreased ASR. Therefore, the better attack strategy is to attack the neutral speech.
@@ V Conclusion
This paper analyzes the differences between backdoor attacks in the domains of images and speech. We proposed EmoAttack, a backdoor attack method based on emotional voice conversion. This method preserves speech’s linguistic content and timbre characteristics while modifying a higher-level attribute of speech: emotion. After EmoAttack training, the emotional utterances can lead the victim model to produce wrong predictions. We conducted backdoor attack experiments on two speech classification tasks. The experimental results demonstrate the excellent attack effectiveness of the EmoAttack. Additionally, we verified that different emotions as target labels result in varying efficiency of the trigger. The intense emotions gain better results. The proposed method aims to provide insights into backdoor attacks in the speech domain.